A company is planning to make an investment of $100,000 in a Machine. The
company’s analyst had estimated that the useful life of the Machine is 5 years and
that each year (from Year 1 to Year 5), the company will receive a net income of
$35,000 from this investment. The acceptable cost of capital is assumed to be 10%
p.a. Calculate:
(a) The payback period of this project/investment plan?
1
Expert's answer
2018-05-11T15:56:26-0400
A company is planning to make an investment of $100,000 in a Machine. The
company’s analyst had estimated that the useful life of the Machine is 5 years and
that each year (from Year 1 to Year 5), the company will receive a net income of
$35,000 from this investment. The acceptable cost of capital is assumed to be 10% p.a. Calculate:
(a) Let's build a table of cumulative cash flow:
year CF Cumulative CF
0 -100000 -100000
1 35000 -65000
2 35000 -30000
3 35000 5000
4 35000 40000
5 35000 75000

The PP = years full recovery + (unrecovered cost at beginning of last year)/(cash flow in following year)=2+30000/35000 = 2.86